OpenStudy (anonymous):

solve for x: 3(3x-4)+5(6x-7)=2x-14

OpenStudy (saifoo.khan):

9x-12 + 30x -35 = 2x - 14 Add like terms: 39x - 57 = 2x - 14 39x - 2x = 57-14 37x = 37 x = 37/37 x =1